OpenStudy (anonymous):

a rectangle has a perimeter 84 cm and length 35 cm. what is its width?

OpenStudy (anonymous):

 Perimeter = 35 + 35 + x + x = 84 70 + 2x = 84

OpenStudy (anonymous):

From what apple said it can be extrapolated that x = 7!